OpenStudy (anonymous):

Slove: 13x-4-9x+20=0 A. -4 B.-12/11 c.-6 d. 6

OpenStudy (igreen):

13x - 4 - 9x + 20 = 0 13x - 9x + 16 = 0 13x - 9x = -16 4x = -16 x = -4

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @ReneeSimone